OpenStudy (anonymous):

if x=2a-b^2 then a equals

OpenStudy (anonymous):

add both sides by \(b^2\), then divide both sides by 2. \(x+b^2 = 2a\) \(\dfrac{x+b^2}{2} = a\)

OpenStudy (anonymous):

Thanks.